OpenStudy (anonymous):

Integral(e^(ab))da= e^(ab)*b^-1 Why?

OpenStudy (anonymous):

As, going the other way: if I differentiate the right equation with respect to a, b doesn't dissapear, as it is a constant, then why does a b^-1 materialise in the antiderivative?

OpenStudy (amistre64):

i take it thats a partial and not b(a)?

OpenStudy (amistre64):

\[D_a(exp({ab}))=exp(ab)*(ab)'\] \[D_a(exp({ab}))=exp(ab)(a'b+ab')\] or maybe if bs a constant \[D_a(exp({ab}))=exp(ab)*b\]

OpenStudy (amistre64):

we need the 1/b to catch it; 1/b * b = 1

OpenStudy (amistre64):

in other words, if we dont put something there to catch the "b" that flies out we dont get the derivative we are looking for

OpenStudy (amistre64):

Dx is a way to notate a derivative operation on something.

OpenStudy (amistre64):

Da means take the derivative of this stuff with respect to a

OpenStudy (amistre64):

go ahead and replace b with your favorite constant thats not 0

OpenStudy (amistre64):

youll see that if we dont have a 1/2 in the antiderivative that we end up with : exp(2a)*2 which is NOT what we are trying to undo.

OpenStudy (amistre64):

so we apply a useful form of 1 into the antiderivative to help us out; since 1* anything doesnt change its value; we need a "2", or a "b" as the case may be, so lets use 2/2 and just pull out the 1/2 for later

OpenStudy (anonymous):

Thanks, after writing that out it clicked.

OpenStudy (amistre64):

:) youre welcome